OpenStudy (anonymous):

Simplify using positive exponents a^4b^-3/ab^-2

OpenStudy (anonymous):

\[\frac{ a^4b ^{-3} }{ ab ^{-2} }\] is the problem writen properly

OpenStudy (anonymous):

\[\frac{a^4}{a}=a^{4-1}=a^3\] is a start

OpenStudy (anonymous):

then \[\frac{b^{-3}}{b^{-4}}=b^{-3-(-4)}=b^{-3+4}=b^1=b\]